OpenStudy (anonymous):

a person 100 feet from the base of a tower measure an angle of 60 degree from the ground to the top of the tower find the height of the tower? (2)what is an angel of measure if the person is 200 feet from the base?

OpenStudy (nowhereman):

Use the definition of the cosine in a right-angled triangle.

OpenStudy (dumbcow):

haha i think i already answered this in a prev post

OpenStudy (dumbcow):

the person makes a right triangle with the base and top of tower height is the opposite side from the angle side 100 ft is adjacent side to angle tan(60) = opposite/adjacent = h/100 h = 100*tan(60) b) tan(x) = h/200 x = inverseTan(h/200), use h found above